# Multi-target quantum compilation algorithm

### v1

@@ 2.2 Applications
Quantum state preparation using quantum algorithms is well-studied for pure states [3, 5, 6]. However, preparing mixed states requires purification first and then preparing the pure state [31]. The conventional purification method needs $2N$ qubits to prepare a mixed state of $N$ qubits. Here, we propose a “dense-purification” method that needs only $N$ qubits.
@@
A quantum state $\rho$ in an $n$-dimensional Hilbert space $\mathscr{H}_{A}$ can be represented using its eigenvalues and eigenstates as $\rho=\sum_{j=1}^{n}p_{j}|j\rangle\langle j|$. In conventional purification, we create a pure state $|\psi\rangle$ in a larger Hilbert space $\mathscr{H}_{A}\otimes\mathscr{H}_{B}$, where $\mathscr{H}_{B}$ is another $n$-dimensional Hilbert space with an orthonormal basis $\{|j\rangle_{B}\}$. The pure state is $|\psi\rangle=\sum_{j}\sqrt{p_{j}}|j\rangle_{A}|j\rangle_{B}$, purifying $\rho$ such that $\rho={\rm Tr}_{B}[|\psi\rangle\langle\psi|]$. In our dense-purification method, we define a pure state $|\psi\rangle=\sum_{j}\sqrt{p_{j}}|j\rangle$, directly representing the dense-purified state of $\rho$. It can be shown that $\rho={\rm diag}(|\psi\rangle\langle\psi|)$, allowing us to extract various properties of $\rho$ from $|\psi\rangle$.
@@
We demonstrate the preparation of thermal equilibrium states, specifically Gibbs states at fixed temperatures. These states are crucial for various applications, such as quantum simulation [32], quantum machine learning [33], quantum condensed matter [34], quantum field theory [35], and cosmology [36, 37]. We focus on the Gibbs state of a transverse field Ising model (TFIM) [38], which is useful for studying thermal phase transitions in condensed matter physics.
@@
Conventionally, to prepare this mixed state on a quantum computer, previous approaches used the conventional purification scheme as [4, 39, 40, 41]
@@
where $Z(\beta)=\sum_{j}e^{-\beta E_{j}}$. Here, $E_{j}$ and $|j\rangle$ are the eigenvalues and eigenstates of $H$, i.e., $H|j\rangle=E_{j}|j\rangle$. The purified state is known as thermofield double (TFD) state [4], which related to the Gibbs state through $\rho(\beta)={\rm Tr}_{B}\big{[}|\psi(\beta)\rangle\langle\psi(\beta)|\big{]}$.
@@
Previously, QAOA [42, 39, 40, 4] and parametrized circuits [41] were used to prepare the quantum state in (7). Here, we apply our GA-VQA to prepare the states in Eqs.(7, 8). We evaluate the closeness of the prepared state to the target state using fidelity and purity metrics
@@
where $\check{\rho}(\beta)={\rm Tr}\big{[}|\psi(\beta)\rangle\langle\psi(\beta)|\big{]}$ for conventional purification and $\check{\rho}(\beta)={\rm diag}\big{[}|\psi_{\rm dense}(\beta)\rangle\langle\psi_{\rm dense}(\beta)|\big{]}$ for dense purification. See the Method section for details on preparing $|\psi(\beta)\rangle$ and $|\psi_{\rm dense}(\beta)\rangle$.
Figure 3: Thermal states preparation. (a) Comparison of fidelity and purity versus $\beta$ at $N=2$ for various methods, alongside theoretical predictions. (b) Quantum circuits used in dense and conventional methods. (c) Fidelity plotted against $N$ at $\beta=2$. (d) Square error of fidelity plotted against $N$ at $\beta=2$. (e) Purity plotted against $N$ at $\beta=2$. (f) Square error of purity plotted against $N$ at $\beta=2$.
In Fig. 3a, we set $N=2$ and examine $\mathcal{F}$ and $\mathcal{P}$ while comparing them in different methods. We demonstrate that both methods align well with the theory, proving the efficacy of the preparation method for creating Gibbs states. In Fig. 3b, we show the corresponding quantum circuits, where the dense method requires fewer qubits and lower circuit depth compared to the conventional method. Next, we focus on the dense method with $\beta=2$. Figure 3c shows the fidelity versus the number of qubits $N$ with the results closely approaching one (theory) for all $N$, and the square error is on the order of $10^{-5}$ as depicted in Fig. 3d. Similarly, we analyze the purity and its error in Fig. 3(e,f), which align well with the theory.
Time-dependent quantum dynamic simulation (TD-QDS)
@@
The local magnetization is shown in Fig. 4, comparing the exact result with various methods, including GA-VQA, Trotterlization [43, 44] with first order (Trotter1) and second order (Trotter2), and Adaptive Variational Quantum Dynamics Simulations (AVQDS) [7]. We set $N=2,J=1,T=10$ and consider several models for $u$ and $h$ as shown in the figure. For the theoretical computation, the local magnetization is given by Eq. (12), with the final state directly computed from $U(t)$ yielding $|\psi(t)\rangle=U(t)|\psi_{0}\rangle$. For the other simulation methods, the local magnetization is derived from the measured probabilities of the final circuit. See detailed calculation in the Method section.
The GA-VQA results closely match the exact results over time for all models, demonstrating the method’s reliability and versatility in capturing quantum dynamics. Trotter2 also achieves high accuracy, aligning well with exact results in cases (a) and (b), but shows deviations in case (c). Trotter1 and AVQDS match well with theoretical predictions only in case (a) and deviate as time $t$ increases in cases (b) and (c). Deviations in Trotter1 are common, as noted in previous studies [43, 45, 44], while discrepancies in AVQDS results stem from simplifications in our calculations, as discussed in the Method section.
3 Discussion
3.1 Demo application for Variational Quantum Eigensolver (VQE)
In this section, we extend the proposed mutil-target compilation approach into VQE. It is a quantum algorithm used to estimate the lowest eigenvalue of a given Hamiltonian, which is crucial in quantum chemistry and optimization problems. It employs a parameterized quantum circuit to prepare trial states and a classical optimization routine to adjust these parameters iteratively, converging towards the lowest eigenvalue. This algorithm is handy in problems where the direct calculation of eigenvalues is computationally expensive, making it a promising candidate for near-term quantum computer applications.
The unitary coupled cluster with singles and doubles (UCCSD) ansatz was used in the first proposal of the VQE algorithm [46]. While achieving high accuracy and attracting significant research interest, the UCCSD ansatz requires a large number of gates and a high depth [47], due to their “staircase” structure. This makes it less suitable for the Noisy Intermediate-Scale Quantum (NISQ) era.
@@ 3.1 Demo application for Variational Quantum Eigensolver (VQE)
$\mathcal{L}_{\rm VQE}(\bm{\Theta},V_{l})=\sum_{j=1}^{n}\langle\textbf{0}|V_{l}^{\dagger}(\bm{\theta}_{j})H_{j}V_{l}(\bm{\theta}_{j})|\textbf{0}\rangle\;;\forall 1\leq l\leq n_{\mathbf{\mathcal{V}}}.$ (13)
where $|\textbf{0}\rangle\equiv|0\rangle^{\otimes N}$ is a reference state, $H_{j}\in\mathcal{H}\ \forall j=\{1,\cdots,n\}$ is a set of $n$ Hamiltonians with respect to the molecular distances $R=\{R_{1},R_{2},\cdots,R_{n}\}$, and $V_{l}$ is a quantum circuit found by GA-VQA scheme.
We apply the GA-VQA scheme to the Hydrogen molecule. In Fig. 5a, we estimate the ground state potential energy surface by optimizing VQE with the best ansatz from the GA-VQA method for ten points ($n=10$), achieving accuracy comparable to the UCCSD ansatz. Here, we approximate the UCCSD ansatz with only the first-order Suzuki-Trotter approximation. Figure 5b shows the convergence of fidelity versus $n_{\rm gene}$ for $d=$ 15, 18, 20, 24, and 25, with their best fidelities plotted in Fig. 5c. As seen in Fig. 5d, the ansatzes identified by GA-VQA possess more parameters compared to UCCSD ansatz but require significantly less depth. Even with the first-order Trotter decomposition, the UCCSD ansatz demands a depth of 74. Consequently, the GA-VQA found ansatzes are more suitable for the NISQ era.
3.2 The complexity of GA-VQA
We discuss here the complexity of GA-VQA. The time complexity of GA-VQA is defined as $\mathcal{O}(n_{\rm gene}n_{\mathcal{V}}\mathcal{F})$ which depends on the number of generations $n_{\rm gene}$, the number of circuits in each generation $n_{\mathcal{V}}$ and the complexity of the fitness function $\mathcal{F}$. When the fitness of circuits in one generation can be evaluated concurrent, the actual time complexity is $\mathcal{O}(\frac{n_{\rm gene}n_{\mathcal{V}}\mathcal{F})}{n_{\rm C}})$ where ${n_{\rm C}}$ is the number of CPU cores. The space complexity is $O(n_{\mathcal{V}}Nd)$, where $N,d$ are the number of qubits and circuit depth, respectively.
4 Method
4.1 Structure of GA
@@
Table 2: List of terminologies used in the common genetic science and theirs counterpart in quantum circuits.
@@ 1 DNA/RNA genetic material quantum gate unitary operator
In a GA scheme, each quantum circuit (individual gene) will be randomly generated from a pool gate with a fixed circuit depth, which consists of various types of quantum gates with one qubit, two-qubits, to one-parameter, and two-parameter. The more gates used, the more possible candidates. This increases the likelihood of finding a suitable candidate but at the same time expands the search space, making us spend more time. Furthermore, using basic gates such as Clifford $\text{Clifford set}=\{{\rm H,S,CX}\}$ [48] enables to implementation of candidates in a real quantum computer, two-qubits gates will be used restrictively. In general, our pool is $\{{\rm H,S,CX,R}_{i}(.)\}$ with $i\in\{x,y,z\}$.
@@
4.2 Selection-Crossover-Mutation
@@
In this work, we used Elitist Selection, One-point Crossover, and Bit Flip Mutation as the default option. We plan to investigate other combinations in future studies. After generating a set of quantum circuits $\mathcal{V}$, we evaluate them using cost functions like fidelity and retain only two best candidates as elitist circuits for the next generation. These selected candidates are paired for crossover to create new ones. Each pair of parents ($p_{1},p_{2}$) is divided into four parts $\{p_{11},p_{12},p_{21},p_{22}\}$ at one point, normally center point. Two new candidates ($c_{1},c_{2}$) are formed by combining these parts: $c_{1}=\{p_{11},p_{22}\}$ and $c_{2}=\{p_{21},p_{12}\}$. In each generation, there is a small probability ($1\%$) that any gate (bit) in a candidate will mutate and be replaced (flip) with a different gate from the pool.
4.3 Detailed experimental setting in the Results
All numerical findings are implemented in Python using Qiskit 0.45.1 with QASMSimulatorPy simulation to verify algorithm convergence. We benchmark $2$ to $10$ qubits on several computer systems, including 24 nodes CPU Intel Xeon X5675 High-performance computer (HPC) at Vietnam Academy Science & Technology, Intel X299-GPU A6000 Workstation and AMD EPYC 7713P cluster @ Tohoku University, and Intel Core i9-10940X CPU at NAIST, Japan.
We mainly use gradient-based optimization with Adam optimizer (except for the VQE), where a set of $\bm{\theta}$ is updated through
$\bm{\theta}^{k+1}=\bm{\theta}^{k}-\alpha\frac{\hat{m}_{k}}{\sqrt{\hat{v}_{k}}+\epsilon},$ (14)
where $m_{k}=\beta_{1}m_{k-1}+\left(1-\beta_{1}\right)\nabla_{\bm{\theta}}\mathcal{C}(\bm{\theta}),v_{k}=\beta_{2}v_{k-1}+(1-\beta_{2})\nabla_{\bm{\theta}}^{2}\mathcal{C}(\bm{\theta}),\hat{m}_{k}=m_{k}/\left(1-\beta_{1}^{k}\right),\hat{v}_{k}=v_{k}/\left(1-\beta_{2}^{k}\right),$ with the hyper-parameters are chosen as $\alpha=0.2,\beta_{1}=0.8,\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The gradient $\partial_{\bm{\theta}}\mathcal{C}(\bm{\theta})$ is given through the general parameter-shift rule [49].
Numerical benchmarking

### v2

@@ III.1 Thermal state preparation (TSP)
Quantum state preparation using quantum algorithms is well-studied for pure states Kuzmin2020variationalquantum ; castro2024variational ; hai2023variational . However, preparing mixed states requires purification first, followed by the preparation of the pure state ezzell2023quantum . The conventional purification method needs $2N$ qubits to prepare a mixed state of $N$ qubits. Here, we propose a “dense-purification” method that needs only $N$ qubits.
@@
A quantum state $\rho$ in a $u$-dimensional Hilbert space $\mathscr{H}_{A}$ can be represented using its eigenvalues and eigenstates as $\rho=\sum_{j=1}^{u}p_{j}|j\rangle\langle j|$. In conventional purification, we create a pure state $|\psi\rangle$ in a larger Hilbert space $\mathscr{H}_{A}\otimes\mathscr{H}_{B}$, where $\mathscr{H}_{B}$ is another $u$-dimensional Hilbert space with an orthonormal basis $\{|j\rangle_{B}\}$. The pure state is $|\psi\rangle=\sum_{j}\sqrt{p_{j}}|j\rangle_{A}|j\rangle_{B}$, purifying $\rho$ such that $\rho={\rm Tr}_{B}[|\psi\rangle\langle\psi|]$. In our dense-purification method, we define a pure state $|\psi\rangle=\sum_{j}\sqrt{p_{j}}|j\rangle$, directly representing the dense-purified state of $\rho$. It can be shown that $\rho={\rm diag}(|\psi\rangle\langle\psi|)$, allowing us to extract various properties of $\rho$ from $|\psi\rangle$.
@@
We demonstrate the preparation of thermal equilibrium states, specifically Gibbs states at fixed temperatures. These states are crucial for various applications, such as quantum simulation 9571985 , quantum machine learning PhysRevX.8.021050 , quantum condensed matter PRXQuantum.4.010305 , quantum field theory Lewin2021 , and cosmology ISRAEL1976107 ; Gao2017 . We focus on the Gibbs state of a transverse field Ising model (TFIM) 10.21468/SciPostPhys.6.3.029 , which is useful for studying thermal phase transitions in condensed matter physics.
Figure 3: Thermal states preparation. (a) Comparison of fidelity and purity versus $\beta$ at $N=2$ for various methods, alongside theoretical predictions. (b) Quantum circuits used in dense and conventional methods. (c) Fidelity plotted against $N$ at $\beta=2$. (d) Square error of fidelity plotted against $N$ at $\beta=2$. (e) Purity plotted against $N$ at $\beta=2$. (f) Square error of purity plotted against $N$ at $\beta=2$. The legends for (c) and (e) are the same as (d) and (f).
@@
Conventionally, to prepare this mixed state on a quantum computer, previous approaches used the conventional purification scheme as Sagastizabal2021 ; PhysRevLett.123.220502 ; Zhu_2020 ; PhysRevApplied.16.054035
@@
where $Z(\beta)=\sum_{j}e^{-\beta E_{j}}$. Here, $E_{j}$ and $|j\rangle$ are the eigenvalues and eigenstates of $H$, i.e., $H|j\rangle=E_{j}|j\rangle$. The purified state is known as thermofield double (TFD) state Sagastizabal2021 , which related to the Gibbs state through $\rho(\beta)={\rm Tr}_{B}\big{[}|\psi(\beta)\rangle\langle\psi(\beta)|\big{]}$.
@@
Previously, QAOA Premaratne_2020 ; PhysRevLett.123.220502 ; Zhu_2020 ; Sagastizabal2021 and parametrized circuits PhysRevApplied.16.054035 were used to prepare the quantum state in (5). Here, we apply our GA-VQA to prepare the states in Eqs. (5, 6). We evaluate the closeness of the prepared state to the target state using fidelity and purity metrics
@@
where $\check{\rho}(\beta)={\rm Tr}\big{[}|\psi(\beta)\rangle\langle\psi(\beta)|\big{]}$ for conventional purification and $\check{\rho}(\beta)={\rm diag}\big{[}|\psi_{\rm dense}(\beta)\rangle\langle\psi_{\rm dense}(\beta)|\big{]}$ for dense purification. See B for details on preparing $|\psi(\beta)\rangle$ and $|\psi_{\rm dense}(\beta)\rangle$.
In Figure  3(a), we set $N=2$ and examine $\mathcal{F}$ and $\mathcal{P}$ while comparing them in different methods. We demonstrate that both methods align well with the theory, proving the efficacy of the preparation method for creating Gibbs states. In Figure  3(b), we show the corresponding quantum circuits, where the dense method requires fewer qubits and lower circuit depth compared to the conventional method. Next, we focus on the dense method with $\beta=2$. Figure 3(c) shows the fidelity versus the number of qubits $N$ with the results closely approaching one (theory) for all $N$, and the square error is on the order of $10^{-5}$ as depicted in Figure  3(d). Similarly, we analyze the purity and its error in Figure  3(e,f), which align well with the theory.
III.2 Time-dependent quantum dynamic simulation (TD-QDS)
@@
The local magnetization is shown in Figure  4, comparing the exact result with various methods, including GA-VQA, Trotterlization Ikeda2023minimum ; ikeda2023trotter24 with first order (Trotter1) and second order (Trotter2), and adaptive variational quantum dynamics simulations (AVQDS) PRXQuantum.2.030307 . We set $N=2,J=1,T=10$ and consider several models for $u$ and $h$ as shown in the figure. For the theoretical computation, the local magnetization is given by Eq. (10), with the final state directly computed from $U(t)$ yielding $|\psi(t)\rangle=U(t)|\psi_{0}\rangle$. For the other simulation methods, the local magnetization is derived from the measured probabilities of the final circuit. See detailed calculation in B.
The GA-VQA results closely match the exact results over time for all models, demonstrating the method’s reliability and versatility in capturing quantum dynamics. Trotter2 also achieves high accuracy, aligning well with exact results in cases (a) and (b), but shows deviations in case (c). Trotter1 and AVQDS match well with theoretical predictions only in case (a) and deviate as time $t$ increases in cases (b) and (c). Deviations in Trotter1 are common, as noted in previous studies Ikeda2023minimum ; zhao2023adaptive ; ikeda2023trotter24 , while discrepancies in AVQDS results stem from simplifications in our calculations, as discussed in B.
III.3 Demo application for Variational Quantum Eigensolver (VQE)
In this section, we extend the proposed multi-target compilation approach into VQE. It is a quantum algorithm used to estimate the lowest eigenvalue of a given Hamiltonian, which is crucial in quantum chemistry and optimization problems. It employs a parameterized quantum circuit to prepare trial states and a classical optimization routine to adjust these parameters iteratively, converging towards the lowest eigenvalue. This algorithm is handy in problems where the direct calculation of eigenvalues is computationally expensive, making it a promising candidate for near-term quantum computer applications.
The unitary coupled cluster with singles and doubles (UCCSD) ansatz was used in the first proposal of the VQE algorithm Peruzzo2014 . While achieving high accuracy and attracting significant research interest, the UCCSD ansatz requires a large number of gates and a high depth Romero_2019 , due to their “staircase” structure. This makes it less suitable for the Noisy Intermediate-Scale Quantum (NISQ) era.
@@ III.3 Demo application for Variational Quantum Eigensolver (VQE)
$\mathcal{L}_{\rm VQE}(\bm{\Theta},V_{l})=\sum_{j=1}^{n}\langle\textbf{0}|V_{l}^{\dagger}(\bm{\theta}_{j})H_{j}V_{l}(\bm{\theta}_{j})|\textbf{0}\rangle\;;\forall 1\leq l\leq n_{\mathbf{\mathcal{V}}},$ (11)
where $|\textbf{0}\rangle\equiv|0\rangle^{\otimes N}$ is a reference state, $H_{j}\in\mathcal{H}\ \forall j=\{1,\cdots,n\}$ is a set of $n$ target Hamiltonians with respect to the molecular distances $R=\{R_{1},R_{2},\cdots,R_{n}\}$, and $V_{l}$ is a quantum circuit found by the GA-VQA scheme.
We apply the GA-VQA scheme to the Hydrogen molecule. In Figure  5(a), we estimate the ground state potential energy surface by optimizing VQE with the best ansatz from the GA-VQA method for ten points ($n=10$), achieving accuracy comparable to the UCCSD ansatz. Here, we approximate the UCCSD ansatz with only the first-order Suzuki-Trotter approximation. Figure 5(b) shows the convergence of fidelity versus $n_{\rm gene}$ for $d=$ 15, 18, 20, 24, and 25, with their best fidelities plotted in Figure  5(c). As seen in Figure  5(d), the ansatzes identified by GA-VQA possess more parameters compared to UCCSD ansatz but require significantly less depth. Even with the first-order Trotter decomposition, the UCCSD ansatz demands a depth of 74. Consequently, the GA-VQA found ansatzes are more suitable for the NISQ era. Details on the technical methods used in these findings are provided in B.
IV Conclusion
We introduced a novel multi-target quantum compilation algorithm aimed at optimizing quantum circuits for multiple objectives simultaneously. By leveraging a genetic algorithm (GA) combined with variational quantum algorithms (VQAs), we were able to efficiently optimize both the structure and parameters of quantum circuits. Our benchmarks and case studies demonstrated the algorithm’s capability to outperform traditional approaches, particularly in scenarios requiring simultaneous optimization, such as simulating thermal states and time-dependent systems. This method offers a significant advancement in quantum compilation, providing a foundation for further exploration in multi-target optimization and its potential applications in quantum computing.
Although this work focuses on developing the multi-target quantum compilation algorithm, analyzing its performance in noisy environments, especially on NISQ devices, is crucial. Noise can affect both the convergence of the genetic algorithm and the fidelity of the compiled circuits. In future works, we will include noise models to assess the algorithm’s robustness and explore strategies to mitigate noise, enhancing convergence and maintaining high fidelity.
Acknowledgements.
L.B.H. thanks Dr. Sahel Ashhab for the fruitful discussions. V.T.H. expresses gratitude to Dr. Luong Ngoc Hoang for advice on genetic algorithms. This research is funded by JSPS KAKENHI Grant Number 23K13025, Unitary Fund, and Tohoku University FRIS URO. L.N.T is partially supported by Vietnam National University Ho Chi Minh City (VNU-HCM) under grant number C2024-28-04.
Data availability
Code availability
The code is available at https://github.com/vutuanhai237/GA-QAS.
Author contributions statement
Competing interests
The author declares no competing interests.
Appendix A Structure of GA
A.1 GA scheme
@@
Table 2: List of terminologies used in the common genetic science and their counterpart in quantum circuits.
@@ 1 DNA/RNA genetic material quantum gate unitary operator
In a GA scheme, each quantum circuit (individual gene) will be randomly generated from a pool gate with a fixed circuit depth, which consists of various types of quantum gates with one qubit, two qubits, to one-parameter, and two-parameter. The more gates used, the more possible candidates. This increases the likelihood of finding a suitable candidate but at the same time expands the search space, making us spend more time. Furthermore, using basic gates such as Clifford $\text{Clifford set}=\{{\rm H,S,CX}\}$ gottesman1998theory enables to implementation of candidates in a real quantum computer, two-qubit gates will be used restrictively. In general, our pool is $\{{\rm H,S,CX,R}_{i}(.)\}$ with $i\in\{x,y,z\}$. Although the H and S gates can be expressed using $R_{i}(.)$, they are essential for reducing the complexity of the ansatz, including lowering the number of parameters and minimizing circuit depth.
@@
A.2 Selection-Crossover-Mutation
@@
In this work, we used Elitist Selection, One-point Crossover, and Bit Flip Mutation as the default option. We plan to investigate other combinations in future studies. After generating a set of quantum circuits $\mathcal{V}$, we evaluate them using cost functions like fidelity and retain only the two best candidates as elitist circuits for the next generation. These selected candidates are paired for crossover to create new ones. Each pair of parents ($p_{1},p_{2}$) is divided into four parts $\{p_{11},p_{12},p_{21},p_{22}\}$ at one point, normally center point. Two new candidates ($c_{1},c_{2}$) are formed by combining these parts: $c_{1}=\{p_{11},p_{22}\}$ and $c_{2}=\{p_{21},p_{12}\}$. In each generation, there is a small probability ($1\%$) that any gate (bit) in a candidate will mutate and be replaced (flip) with a different gate from the pool.
Appendix B Detailed experimental setting
All numerical findings are implemented in Python using Qiskit 0.45.1 with QASMSimulatorPy simulation to verify algorithm convergence. We benchmark $2$ to $10$ qubits on several computer systems, including 24 nodes CPU Intel Xeon X5675 High-performance computer (HPC) at Vietnam Academy Science & Technology, Intel X299-GPU A6000 Workstation and AMD EPYC 7713P cluster at Tohoku University, and Intel Core i9-10940X CPU at NAIST, Japan.
We mainly use gradient-based method with Adam optimizer (except for the VQE), where a set of $\bm{\theta}$ is updated through
$\bm{\theta}^{k+1}=\bm{\theta}^{k}-\alpha\frac{\hat{m}_{k}}{\sqrt{\hat{v}_{k}}+\epsilon},$ (B.1)
where $m_{k}=\beta_{1}m_{k-1}+\left(1-\beta_{1}\right)\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}),v_{k}=\beta_{2}v_{k-1}+(1-\beta_{2})\nabla_{\bm{\theta}}^{2}\mathcal{L}(\bm{\theta}),\hat{m}_{k}=m_{k}/\left(1-\beta_{1}^{k}\right),\hat{v}_{k}=v_{k}/\left(1-\beta_{2}^{k}\right),$ with the hyper-parameters are chosen as $\alpha=0.2,\beta_{1}=0.8,\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The gradient $\partial_{\bm{\theta}}\mathcal{L}(\bm{\theta})$ is given through the general parameter-shift rule Hai2024 .
B.1 Numerical benchmarking